A Bibliography of Supercomputing ’89

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

07 August 2013 Version 1.04

Abstract

This bibliography records articles presented at the Supercomputing ’89 conference.

Title word cross-reference

2 [MA89].

/MP [MA89].

10 [Fat89]. 10-Q [Fat89]. 1024-processor [GBSS89].

2 [Fat89]. 29-state [Sig89].

3/SX [WMT89]. 3090 [AG89a].

64-bit [KM89].

’89 [ACM89].

academic [MKYK89]. acoustics [LSS89]. activation [YHKM89]. administration [BAC^{+}89]. aerothermodynamics [Dei89]. AHEAD [dEV89]. algorithm [Bis89a, LC89, PC89, ZWM89]. algorithms [AG89a, CP89, GD89b, Iye89, OS89, PBP89, Sha89]. alignment [Iye89]. allocation [JCC89]. among [Cha89]. analysis [Cal89, EGP89, Fat89, HK89, MJ89]. analytical [AG89b]. approach [JCC89]. architectural [BCS^{+}89]. architecture [SMCG89, WMT89]. architectures [And89, LC89, Mv89]. area [Lek89]. array [GD89b, SHO^{+}89]. assembly [WW89]. assignment [BNR89, PC89]. assisted [Par89]. associated [PRWB89]. associative [PWY^{+}89]. atmospheric [JH89]. Automatic [BNR89, TKA89]. Automatically [NE89]. automaton [Sig89]. balance [WL89]. balancing [CP89]. based [GD89a, Hei89]. Benchmark [CHK89, BCS^{+}89]. beyond [DSOZ89]. bit [KM89]. block [Bak89, Bis89a]. blocking [Hei89]. bound [Par89]. Building [AG89b].

C [MA89]. cache [MA89]. cache-coherent [MA89]. calculations [CHK89]. can [Sig89]. capabilities [Kut89]. Capability [FT89]. Carlo [BCS^{+}89, WW89]. case [Sig89]. cellular [Sig89]. channel [OJ89]. character [TKA89]. characteristics [Iye89]. Chare [SK89]. Chare-Kernel [SK89]. checksums [Che89]. chemistry [FHS89]. Cholesky [AG89a]. chromodynamics [SHO^{+}89]. circuit [FYH89, Hei89, PBP89]. circuit-switched [Hei89]. circulant [Bak89]. code [Car89]. code-methods [Car89]. coherent [MA89]. Combining [RJ89]. Communication [Gro89, SR89]. communities [MKYK89]. Comparative [Hei89]. Comparison [NE89]. Compilers [NE89]. complex [RS89, Sig89]. composition [CT89]. computation [CC89]. Computational [Dei89, Kut89, FT89, LSS89]. Computations [GML89, Mv89]. computer [CHK89, Par89, PWY^{+}89, ZRV89]. computers [HK89, Hua89]. Computing [BW89]. concepts [Cro89]. concerns [Car89]. concurrency [Cha89]. Concurrent [WW89, CT89]. conflict [Cal89]. conflicts [TM89]. conjugate [MJ89]. connected [CC89]. Connection [Dag89, ESH89, JHM89, JKFM89, BBGJ89]. Construction [MKYK89]. constructs [Mv89]. consume [RJ89]. convergence [MJ89]. cost [BW89]. coupled [Mv89]. Cray [Bak89, Fat89, KC89]. Cray-2 [Bak89, Fat89]. crossbar [HT89]. cube [CC89, JCC89]. cubes [GD89a]. current [FT89, Kut89].

d [HJ89]. DASP [PWY^{+}89]. data [BNR89, MJ89, ZWM89]. data-level [ZWM89]. data/program [BNR89]. debuggers [PU89]. debugging [Che89, EGP89]. decomposition [Bak89, CP89, CC89]. depth [Par89]. digital [Hua89]. Dilation [HJ89]. dimensional [OJ89]. directions [Kut89]. dissipation [OJ89]. distributed [BPA89, MA89, PWY^{+}89, SR89]. distributed-memory [BPA89]. due [GML89]. dynamic [SK89]. dynamical [BBGJ89]. dynamics [FT89, Kut89].

editor [BKK^{+}89]. Effective [KC89]. effects [HD89]. Efficient [CC89, PBP89]. Element [MJ89]. elimination [Lev89]. embedding [HJ89]. environment [Bis89b, GGJ89, NNIL89]. equations [OJ89]. ETA [Fat89]. Euler [OJ89]. even [Lev89]. Event [EGP89]. evolution [WW89]. examples [FHS89]. experimentation [GGJ89]. external [Bai89].

factorization [AG89a, Bis89a, SR89].

fermions [BBGJ89]. FFT [JKFM89]. FFTs [Bai89]. File [Cro89, NNIL89]. flow [CHK89, Dag89, JH89, OJ89]. fluid [FT89, Kut89]. fluid-dynamics-current [Kut89]. FORTRAN [dEV89, KKK89, NE89]. framework [HT89]. function [BW89]. future [Kut89]. Futures [Cha89].

’GAMTEB [BCS^{+}89]. general [PWY^{+}89]. general-purpose [PWY^{+}89]. generalized [Lev89]. geometric [OS89]. global [JH89]. gradient [MJ89]. graph [YHKM89]. grid [CHK89]. ground [GML89]. GTS [ALTB89].

Hardware [WMT89]. Harray [YHKM89]. HCN [GD89a]. HDB [Che89]. hierarchical [Bai89]. High [And89, NNIL89, Che89, HK89, SMCG89]. high-speed [HK89]. HVDC [GML89]. hyper [HJ89]. hyper-pyramid [HJ89]. hypercube [GBSS89, HJ89]. hypersonic [Dag89].

I/O [Cro89]. i860 [KM89]. IBM [AG89a]. implement [Sig89]. Implementation [Dag89, CP89]. Implementing [GD89b]. increasing [WW89]. inputs [Par89]. integrated [CP89, HT89]. interactive [AG89b, BKK^{+}89]. interconnection [GD89a]. internet [MKYK89]. inverse [LC89]. issues [Car89]. iteration [Wol89]. iWarp [Gro89].

Japanese [MKYK89].

K9 [BPA89]. Kernel [SK89]. kinematics [LC89]. Knowledge [ESH89].

language [Mv89]. LANL [BCS^{+}89]. Large [dEV89, Car89]. level [Che89, ZWM89, Zia89]. Limits [BAC^{+}89]. linear [GD89b]. Load [CP89]. logic [PBP89]. loop [WL89]. loosely [Mv89]. lower [Par89].

Machine [Dag89, ESH89, JHM89, JKFM89, CC89, Sig89, TM89, BBGJ89]. Macrotasking [Bak89]. manipulations [TKA89]. mapping [Zia89]. matrices [Bak89]. Matrix [JHM89, SR89]. Measuring [ZRV89]. mechanism [Cha89]. Memory [TM89, Bai89, BPA89, Cal89, MA89, RS89, RJ89, WWS89]. mesh [SR89]. method [MJ89]. methodology [RS89]. methods [Car89]. microprocessor [KM89]. migration [JCC89]. MIMD [JH89, PWY^{+}89, SHO^{+}89]. Models [PU89, AG89b, JH89]. Monte [BCS^{+}89, WW89]. MP [Fat89, MA89]. multi [Zia89]. multi-level [Zia89]. multicomputers [RS89]. Multiple [HT89]. multiplication [JHM89]. Multiprocessor [MHHG89, JCC89, MA89, RJ89]. multiprocessor-system [MA89]. multiprocessors [GGJ89, RS89, WWS89].

n [CJ89, JCC89]. N-cube [JCC89]. NEC [WMT89]. NERV [MHHG89]. network [GD89a, HT89, WWS89]. networking [Lek89]. networks [Hei89, MHHG89, Par89]. Neumann [Sig89]. Neural [WWS89, MHHG89]. Nevada [ACM89]. nine [Par89]. November [ACM89]. number [WW89]. numerical [SHO^{+}89].

O [Cro89]. objects [Cha89]. ocean [LSS89]. odd [Lev89]. odd-even [Lev89]. operations [RJ89, TKA89]. operators [OJ89]. Optical [Hua89]. Optimal [Fre89, Par89]. order [MJ89].

PAGE [dEV89]. PAGE-AHEAD [dEV89]. Parallel [Iye89, JH89, LC89, Lev89, Mv89, OS89, Sha89, WL89, AG89a, And89, BKK^{+}89, BPA89, CP89, Cro89, EGP89, FHS89, GD89b, MJ89, PU89, PC89, PWY^{+}89, PBP89, PRWB89, SMCG89, YHKM89, ZWM89, ZRV89]. parallelization [ALTB89]. Parallelizing [Car89, RS89]. ParaScope [BKK^{+}89]. particle [BCS^{+}89, Dag89]. particles [WW89]. partitioning [BNR89, PBP89]. Pascal [TKA89]. peak [CJ89]. pentadiagonal [Lev89]. Performance [HK89, And89, AG89b, BW89, CJ89, Fat89, GGJ89, Hei89, NNIL89, Sha89, SMCG89, TM89]. physics [FHS89, WBP^{+}89]. pipelined [RJ89]. pivoting [Bis89a]. Practical [FHS89]. preceding [YHKM89]. prediction [AG89b, ZWM89]. principle [BNR89]. probability [Hei89]. problem [PC89, Zia89]. problems [OS89]. Proceedings [ACM89]. processing [PWY^{+}89, PRWB89, SMCG89, YHKM89]. processor [GBSS89, JCC89, SR89, WL89]. processors [BW89, BPA89, SHO^{+}89]. produce [RJ89]. product [Mv89]. program [BNR89, GBSS89]. programming [BKK^{+}89, JH89]. Programs [dEV89, CT89, EGP89, RS89]. Protein [ZWM89]. purpose [PWY^{+}89]. pyramid [HJ89].

Q [Fat89]. QCD [BBGJ89]. QCDPAX [SHO^{+}89]. QR [Bis89a]. quadratic [PC89]. quantity [BW89]. quantum [SHO^{+}89].

radar [GBSS89]. radix [JKFM89]. radix-2 [JKFM89]. rank [CJ89]. rank-n [CJ89]. rarefied [Dag89]. rate [MJ89]. Realities [PRWB89]. reconfigurable [SMCG89]. recurrences [ALTB89]. reduction [SR89]. Refined [KKK89]. relational [TKA89]. relativistic [HD89]. Reno [ACM89]. representation [ESH89]. resource [BAC^{+}89]. restricted [Bis89a]. results [Cal89]. return [GML89]. rise [GML89]. robot [LC89].

scalability [ZRV89]. scheduling [SK89]. scheme [YHKM89]. scientific [Car89]. searching [OS89]. security [Bis89b]. selection [Fre89]. sequence [Iye89]. service [Lek89]. shared [RJ89, WWS89]. shared-memory [WWS89]. SIMD [CC89, CHK89, Sig89]. simulation [Dag89, FYH89, GBSS89, MHHG89, PBP89, SHO^{+}89, WWS89, WW89]. simulations [WBP^{+}89]. simulator [BPA89]. single [BNR89]. singular [Bak89, CC89]. soil [GML89]. solution [Lev89]. solver [CJ89, CHK89]. Some [Cal89]. sorting [GD89b, Par89]. space [HD89, Wol89]. sparse [SR89]. speed [BW89, HK89]. spreading [WL89]. state [Sig89]. Static [DSOZ89]. strategy [SK89]. stress [MJ89]. string [TKA89]. structure [ZWM89]. Study [NE89, BCS^{+}89, OJ89, Sig89]. super [Sha89]. supercomputer [WMT89, WBP^{+}89]. Supercomputers [LSS89, Fat89, FT89, KC89]. Supercomputing [ACM89, FYH89, Bis89b, FHS89, HT89, KM89, NNIL89]. superconcurrency [Fre89]. switched [Hei89]. SX [WMT89]. SX-3 [WMT89]. SX-3/SX-X [WMT89]. SX-X [WMT89]. synchronization [DSOZ89, EGP89]. system [BAC^{+}89, MA89, NNIL89, SK89, WMT89, YHKM89]. system-Harray [YHKM89]. systems [CP89, Gro89, Lev89, Zia89, ZRV89].

task [CP89, JCC89]. techniques [CP89]. technology [WMT89]. temperature [GML89]. tensor [Mv89]. theory [Fre89]. three [Fat89, OJ89]. three-dimensional [OJ89]. tight [ALTB89]. tiling [Wol89]. time [HD89]. tool [AG89b, BKK^{+}89, Che89]. transport [BCS^{+}89]. transputers [GD89b]. Tuning [CJ89]. Type [Lek89].

unfolding [YHKM89]. UNIX

[BAC^{+}89, Bis89b]. unstructured [CHK89]. Update [KKK89, CJ89]. use [KC89]. Using [dEV89, BNR89, Bis89a, BCS^{+}89, Che89, Lev89, PWY^{+}89].

value [Bak89, CC89]. Vector [AG89a, Fat89, SHO^{+}89, WWS89]. Vectorization [BCS^{+}89, ALTB89, TKA89]. vectorized [WW89]. Vectorizing [NE89]. versatile [GD89a]. vision [CP89]. Visualization [WBP^{+}89, PU89]. Visualizing [HD89]. VLIW [DSOZ89]. VLSI [LC89].

wavefront [CJ89]. wide [Lek89].

X [WMT89].

Y-MP [Fat89].

References

ACM:1989:PSN

[ACM89] ACM, editor. Proceedings, Supercomputing ’89: November 13–17, 1989, Reno, Nevada. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Agarwal:1989:VPA

[AG89a] R. C. Agarwal and F. G. Gustavson. Vector and parallel algorithms for Cholesky factorization on IBM 3090. In ACM [ACM89], pages 225–233. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Ataputtu:1989:BAM

[AG89b] D. Ataputtu and D. Gannon. Building analytical models into an interactive performance prediction tool. In ACM [ACM89], pages 521–530. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Ayguade:1989:GPV

[ALTB89] E. Ayguade, J. Labarta, J. Torres, and P. Borensztejn. GTS: parallelization and vectorization of tight recurrences. In ACM [ACM89], pages 531–539. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Anderson:1989:HPP

[And89] R. E. Anderson. High performance parallel architectures. In ACM [ACM89], pages 410–415. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Bettison:1989:LSU

[BAC^{+}89] A. Bettison, F. Adcock, P. Chubb, A. Gollan, C. Maltby, and N. Russell. Limits — a system for UNIX resource administration. In ACM [ACM89], pages 686–692. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Bailey:1989:FEH

[Bai89] D. H. Bailey. FFTs in external or hierarchical memory. In ACM [ACM89], pages 234–242. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Baker:1989:MSV

[Bak89] J. R. Baker. Macrotasking the singular value decomposition of block circulant matrices on the cray-2. In ACM [ACM89], pages 243–247. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Baillie:1989:QDF

[BBGJ89] C. F. Baillie, R. G. Brickner, R. Gupta, and L. Johnsson. QCD with dynamical fermions on the Connection Machine. In ACM [ACM89], pages 2–9. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Burns:1989:VMC

[BCS^{+}89] P. J. Burns, M. Christon, R. Schweitzer, O. M. Lubeck, H. J. Wasserman, M. L. Simmons, and D. V. Pryor. Vectorization of Monte Carlo particle transport: an architectural study using the LANL benchmark ’GAMTEB’. In ACM [ACM89], pages 10–20. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Bischof:1989:BQF

[Bis89a] C. H. Bischof. A block QR factorization algorithm using restricted pivoting. In ACM [ACM89], pages 248– 256. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Bishop:1989:USS

[Bis89b] M. Bishop. UNIX security in a supercomputing environment. In ACM [ACM89], pages 693–698. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Balasundaram:1989:PEI

[BKK^{+}89] V. Balasundaram, K. Kennedy, U. Kremer, K. McKinley, and J. Subhlok. The ParaScope editor: an interactive parallel programming tool. In ACM [ACM89], pages 540– 550. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Bic:1989:ADP

[BNR89] L. Bic, M. D. Nagel, and J. M. A.

Roy. Automatic data/program partitioning using the single assignment principle. In ACM [ACM89], pages 551–556. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Beadle:1989:KSD

[BPA89] P. Beadle, C. Pommerell, and M. Annaratone. K9: a simulator of distributed-memory parallel processors. In ACM [ACM89], pages 765– 774. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-

Barton:1989:CPF

[BW89] M. L. Barton and G. R. Withers. Computing performance as a function of the speed, quantity, and cost of the processors. In ACM [ACM89], pages 759–764. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Calahan:1989:SRM

[Cal89] D. A. Calahan. Some results in memory conflict analysis. In ACM [ACM89], pages 775–778. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Carmona:1989:PLS

[Car89] E. A. Carmona. Parallelizing a large scientific code-methods, issues, and concerns. In ACM [ACM89], pages 21–31. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Chuang:1989:ECS

[CC89] H. Y. H. Chuang and L. Chen. Efficient computation of the singular value decomposition on cube connected SIMD machine. In ACM [ACM89], pages 276–282. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Chatterjee:1989:FMC

[Cha89] A. Chatterjee. Futures: a mechanism for concurrency among objects. In ACM [ACM89], pages 562– 567. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Cheng:1989:HHL

[Che89] D. Y. Cheng. HDB — a high level debugging tool using checksums. In ACM [ACM89], pages 568– 573. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Clary:1989:BCU

[CHK89] J. S. Clary, G. A. Howell, and S. L.

Karman, Jr. Benchmark calculations with an unstructured grid flow solver on a SIMD computer. In ACM [ACM89], pages 32–41. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Chang:1989:TRU

[CJ89] S. Chang and M. A. Johnson. Tuning the rank-n update in a wavefront solver for peak performance. In ACM [ACM89], pages 257–265. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Choudhary:1989:LBT

[CP89] A. N. Choudhary and J. H. Patel.

Load balancing and task decomposition techniques for parallel implementation of integrated vision systems algorithms. In ACM [ACM89], pages 266–275. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Crockett:1989:FCP

[Cro89] T. W. Crockett. File concepts for parallel I/O. In ACM [ACM89], pages 574–579. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Chandy:1989:CCP

[CT89] K. M. Chandy and S. Taylor.

The composition of concurrent programs. In ACM [ACM89], pages 557–561. ISBN 0-89791-341-8.

LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Dagum:1989:IHR

[Dag89] L. Dagum. Implementation of a hypersonic rarefied flow particle simulation on the connection machine. In ACM [ACM89], pages 42–50. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Deiwart:1989:CA

[Dei89] G. S. Deiwart. Computational aerothermodynamics. In ACM [ACM89], pages 51–57. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

diAntonio:1989:UPA

[dEV89] R. di Antonio, J. Eilert, and M. Vitaletti. Using PAGE-AHEAD for large FORTRAN programs. In ACM [ACM89], pages 511–520. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Dietz:1989:SSB

[DSOZ89] H. Dietz, T. Schwederski, M. O’Keefe, and A. Zaafrani. Static synchronization beyond VLIW. In ACM [ACM89], pages 416–425. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Emrath:1989:ESA

[EGP89] P. A. Emrath, S. Ghosh, and D. A. Padua. Event synchronization analysis for debugging parallel programs. In ACM [ACM89], pages 580–588. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Evett:1989:KRC

[ESH89] M. Evett, L. Spector, and J. Hendler.

Knowledge representation on the connection machine. In ACM [ACM89], pages 283–293. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Fatoohi:1989:VPA

[Fat89] R. A. Fatoohi. Vector performance analysis of three supercomputers: Cray-2, Cray Y-MP, and ETA 10-Q. In ACM [ACM89], pages 779– 788. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Fox:1989:PPS

[FHS89] G. C. Fox, P. Hipes, and J. Salmon.

Practical parallel supercomputing: examples from chemistry and physics. In ACM [ACM89], pages 58–70. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Freund:1989:OST

[Fre89] R. F. Freund. Optimal selection theory for superconcurrency. In ACM [ACM89], pages 699–703. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Fuji:1989:CCS

[FT89] K. Fuji and Y. Tamura. Capability of current supercomputers for the computational fluid dynamics. In ACM [ACM89], pages 71–80. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Fukui:1989:SCS

[FYH89] Y. Fukui, H. Yoshida, and S. Higono. Supercomputing of circuit simulation. In ACM [ACM89], pages 81–85. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Gustafson:1989:RSP

[GBSS89] J. L. Gustafson, R. E. Benner, M. P. Sears, and T. D. Sullivan. A radar simulation program for a 1024-processor hypercube. In ACM [ACM89], pages 96–105. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Ghose:1989:HVI

[GD89a] K. Ghose and K. R. Desai. The HCN: a versatile interconnection network based on cubes. In ACM [ACM89], pages 426–435. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Goodwin:1989:IPS

[GD89b] W. P. Goodwin, IV and S. K. Das. Implementing parallel sorting algorithms on a linear array of transputers. In ACM [ACM89], pages 789– 796. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Gaur:1989:EPE

[GGJ89] Y. Gaur, V. A. Guarna, Jr., and D. Jablonowski. An environment for performance experimentation on multiprocessors. In ACM [ACM89], pages 589–596. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Greiss:1989:CST

[GML89] H. Greiss, D. Mukhedkar, and P. J. Lagace. Computations of soil temperature rise due to HVDC ground return. In ACM [ACM89], pages 86– 95. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Gross:1989:CIS

[Gro89] T. Gross. Communication in iWarp systems. In ACM [ACM89], pages 436–445. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Hsiung:1989:VRE

[HD89] P.-K. Hsiung and R. H. P. Dunn. Visualizing relativistic effects in space time. In ACM [ACM89], pages 597– 606. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Hein:1989:CPC

[Hei89] C. Hein. Comparative performance of circuit-switched networks based on blocking probability. In ACM [ACM89], pages 704–712. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Ho:1989:DDE

[HJ89] C.-T. Ho and S. L. Johnsson. Dilation d embedding of a hyperpyramid into a hypercube. In ACM [ACM89], pages 294–303. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Helin:1989:PAH

[HK89] J. Helin and K. Kaski. Performance analysis of high-speed computers. In ACM [ACM89], pages 797–808. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Hoebelheinrich:1989:MCN

[HT89] R. Hoebelheinrich and R. Thomsen. Multiple crossbar network integrated supercomputing framework. In ACM [ACM89], pages 713–720. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Huang:1989:ODC

[Hua89] A. Huang. Optical digital computers. In ACM [ACM89], pages 446– 449. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Iyengar:1989:PCS

[Iye89] A. K. Iyengar. Parallel characteristics of sequence alignment algorithms. In ACM [ACM89], pages 304–313. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Jang:1989:NAP

[JCC89] J. E. Jang, S. W. Choi, and W. K. Cho. A new approach to processor allocation and task migration in an N-cube multiprocessor. In ACM [ACM89], pages 314–325. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Jakob:1989:PMP

[JH89] R. Jakob and J. J. Hack. Parallel MIMD programming for global models of atmospheric flow. In ACM [ACM89], pages 106–112. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Johnsson:1989:MMC

[JHM89] S. L. Johnsson, T. Harris, and K. K. Mathur. Matrix multiplication on the connection machine. In ACM [ACM89], pages 326–332. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Johnsson:1989:RFC

[JKFM89] S. L. Johnsson, R. L. Krawitz, R. Frye, and D. MacDonald. A radix-2 FFT on the connection machine. In ACM [ACM89], pages 809– 819. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Kramer:1989:EUC

[KC89] W. T. C. Kramer and J. M. Craw.

Effective use of Cray supercomputers. In ACM [ACM89], pages 721– 731. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Klappholz:1989:RFU

[KKK89] D. Klappholz, X. Kong, and A. D.

Kallis. Refined Fortran: An update. In ACM [ACM89], pages 607– 615. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Kohn:1989:ISM

[KM89] L. Kohn and N. Margulis. The i860 64-bit supercomputing microprocessor. In ACM [ACM89], pages 450– 456. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Kutler:1989:CFC

[Kut89] P. Kutler. Computational fluiddynamics-current capabilities and directions for the future. In ACM [ACM89], pages 113–122. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Lai:1989:PAV

[LC89] J. Z. C. Lai and Ming Chao. Parallel algorithm and VLSI architectures for a robot’s inverse kinematics. In ACM [ACM89], pages 123– 132. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Lekashman:1989:TSW

[Lek89] J. Lekashman. Type of service wide area networking. In ACM [ACM89], pages 732–736. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Levit:1989:PSP

[Lev89] C. Levit. Parallel solution of pentadiagonal systems using generalized odd-even elimination. In ACM [ACM89], pages 333–336. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Lee:1989:SCO

[LSS89] D. Lee, M. H. Schultz, and F. Saied. Supercomputers in computational ocean acoustics. In ACM [ACM89], pages 133–140. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Marquardt:1989:CSM

[MA89] D. E. Marquardt and H. S. AlKhatib. C^{2} /MP: a cachecoherent, distributed memory multiprocessor-system. In ACM [ACM89], pages 466–475. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Manner:1989:MSN

[MHHG89] R. Manner, H. Horner, R. Hauser, and A. Genthner. Multiprocessor simulation of neural networks with NERV. In ACM [ACM89], pages 457–465. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Mathur:1989:EOC

[MJ89] K. K. Mathur and S. L. Johnsson. Element order and convergence rate of the conjugate gradient method for data parallel stress analysis. In ACM [ACM89], pages 337– 343. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Murai:1989:CIJ

[MKYK89] J. Murai, H. Kusumoto, S. Yamaguchi, and A. Kato. Construction of internet for Japanese academic communities. In ACM [ACM89], pages 737–746. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Mehrotra:1989:PLC

[Mv89] P. Mehrotra and J. van Rosendale.

Parallel language constructs for tensor product computations on loosely coupled architectures. In ACM [ACM89], pages 616–626. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Nobayashi:1989:CSA

[NE89] H. Nobayashi and C. Eoyang. A comparison study of automatically vectorizing Fortran compilers. In ACM [ACM89], pages 820–825. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Nishino:1989:HPF

[NNIL89] H. Nishino, S. Naka, K. Ikumi, and W. Leslie. High performance file system for supercomputing environment. In ACM [ACM89], pages 747– 756. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Olsson:1989:SDO

[OJ89] P. Olsson and S. L. Johnsson. A study of dissipation operators for the Euler equations and a threedimensional channel flow. In ACM [ACM89], pages 141–151. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Oh:1989:PAG

[OS89] S.-J. Oh and M. Suk. Parallel algorithms for geometric searching problems. In ACM [ACM89], pages 344– 350. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Parberry:1989:CAO

[Par89] I. Parberry. A computer assisted optimal depth lower bound for sorting networks with nine inputs. In ACM [ACM89], pages 152–161. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Patil:1989:ECP

[PBP89] S. Patil, P. Banerjee, and C. D. Polychronopoulos. Efficient circuit partitioning algorithms for parallel logic simulation. In ACM [ACM89], pages 361–370. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Pardalos:1989:PAQ

[PC89] P. M. Pardalos and J. V. Crouse. A parallel algorithm for the quadratic assignment problem. In ACM [ACM89], pages 351–360. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Platt:1989:RAP

[PRWB89] D. E. Platt, A. R. Rossi, J. W.

Wells, and J. Becker. Realities associated with parallel processing. In ACM [ACM89], pages 162–174. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Pancake:1989:MVP

[PU89] C. M. Pancake and S. Utter. Models for visualization in parallel debuggers. In ACM [ACM89], pages 627– 636. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Park:1989:DGM

[PWY^{+}89] Y. K. Park, C. Walter, H. Yee, T. Roden, and S. Berkovich. DASP: a general-purpose MIMD parallel computer using distributed associative processing. In ACM [ACM89], pages 476–484. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Rodriguez:1989:CPC

[RJ89] B. J. Rodriguez and H. F. Jordan.

Combining produce and consume operations in a pipelined shared memory multiprocessor. In ACM [ACM89], pages 485–495. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Ramunujam:1989:MPP

[RS89] J. Ramunujam and P. Sadayappan. A methodology for parallelizing programs for multicomputers and complex memory multiprocessors. In ACM [ACM89], pages 637– 646. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Shakshober:1989:PAS

[Sha89] D. J. Shakshober. Parallel algorithms for super performance. In ACM [ACM89], pages 380–388. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Shirakawa:1989:QMA

[SHO^{+}89] T. Shirakawa, T. Hoshino, Y. Oyanagi, Y. Iwasaki, T. Yoshie, K. Kanaya, S. Ichii, and T. Kawai. QCDPAX — an MIMD array of vector processors for the numerical simulation of quantum chromodynamics. In ACM [ACM89], pages 495– 504. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Signorini:1989:HSM

[Sig89] J. Signorini. How a SIMD machine can implement a complex cellular automaton? a case study: von Neumann’s 29-state cellular automaton. In ACM [ACM89], pages 175–186. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Shu:1989:DSS

[SK89] W. Shu and L. V. Kale. A dynamic scheduling strategy for the charekernel system. In ACM [ACM89], pages 389–398. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Shively:1989:HPR

[SMCG89] R. R. Shively, E. B. Morgan, T. W. Copley, and A. L. Gorin. A high performance reconfigurable parallel processing architecture. In ACM [ACM89], pages 505–509. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Sadayappan:1989:CRD

[SR89] P. Sadayappan and S. K. Rao.

Communication reduction for distributed sparse matrix factorization on a processor mesh. In ACM [ACM89], pages 371–379. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Tsuda:1989:AVC

[TKA89] T. Tsuda, Y. Kunieda, and P. Atipas. Automatic vectorization of character string manipulations and relational operations in pascal. In ACM [ACM89], pages 187–196. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Tang:1989:MCM

[TM89] P. Tang and R. H. Mendez. Memory conflicts and machine performance. In ACM [ACM89], pages 826–831. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Watson:1989:VSS

[WBP^{+}89] V. Watson, G. Bancroft, T. Plessel, F. Merritt, Walatka, and P. P. Visualization of supercomputer simulations in physics. In ACM [ACM89], pages 647–654. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Wu:1989:PPB

[WL89] Y. Wu and T. Lewis. Parallel processor balance through loop spreading. In ACM [ACM89], pages 665– 674. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Watanabe:1989:HTA

[WMT89] T. Watanabe, H. Matsumoto, and P. D. Tannenbaum. Hardware technology and architecture of the NEC SX-3/SX-X supercomputer system. In ACM [ACM89], pages 842–846. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Wolfe:1989:MIS

[Wol89] M. Wolfe. More iteration space tiling. In ACM [ACM89], pages 655–664. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Wu:1989:CVM

[WW89] C.-H. Wu and C.-J. Wang. Concurrent and vectorized Monte Carlo simulation of the evolution of an assembly of particles increasing in number. In ACM [ACM89], pages 205–214. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Wang:1989:NNS

[WWS89] C.-J. Wang, C.-H. Wu, and S. Sivasundaram. Neural network simulation on shared-memory vector multiprocessors. In ACM [ACM89], pages 197–204. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Yamana:1989:PAS

[YHKM89] H. Yamana, T. Hagiwara, J. Kohdate, and Y. Muraoka. A preceding activation scheme with graph unfolding for the parallel processing system-harray. In ACM [ACM89], pages 675–684. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Ziavras:1989:MPM

[Zia89] S. G. Ziavras. On the mapping problem for multi-level systems. In ACM [ACM89], pages 399–408. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Zorbas:1989:MSP

[ZRV89] J. R. Zorbas, D. J. Reble, and R. E.

VanKooten. Measuring the scalability of parallel computer systems. In ACM [ACM89], pages 832–841. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Zhang:1989:PSP

[ZWM89] X. Zhang, D. Waltz, and J. P.

Mesirov. Protein structure prediction by a data-level parallel algorithm. In ACM [ACM89], pages 215–223. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.